OpenStudy (anonymous):

The function f(x)=x^3-3x^2+2x rises as x grows very large. True or false?

OpenStudy (anonymous):

Have you learned derivatives?

OpenStudy (anonymous):

Kind of? but i'm not really understanding what the question is asking

OpenStudy (anonymous):

Yep moreover this question seems a bit vague nobody asks when x gets really larger :p

OpenStudy (anonymous):

Very large*

OpenStudy (anonymous):

So maybe they're asking about the lim of x as it temds to infinity

OpenStudy (anonymous):

actually, a lot of people ask about what happens as x gets 'very large' -- that's the entire point of asymptotics. these are limits \(x\to\pm\infty\) -- you have encountered them as horizontal asymptotes or 'end behavior'. it is a common topic in precalculus treatments of polynomial and rational functions

OpenStudy (anonymous):

Are you familiar with these?

OpenStudy (anonymous):

Well my friend in my math we have no very large :) either we've got limx-->+inf of -inf :)

OpenStudy (anonymous):

so false?

OpenStudy (anonymous):

@mishtea the point is that, for sufficiently large \(x\), \(x^3\) will grow faster and faster than any terms of lower powers of \(x\), like \(x^2,x,1\). so notice that if we factor out \(x^3\) we get $$x^3-3x^2+2x=x^3(1-3/x+2/x^2)$$ you can immediately see that if we let \(x\) grow and grow larger and larger, \(3/x\) and \(2/x^2\) would quickly become negligibly small, so the whole function would behave a lot like $$x^3-3x^2+2x=x^3(1-3/x+2/x^2)\sim x^3(1)=x^3$$ where \(\sim\) means "behaves like for large \(x\)" does \(x^3\) 'rise' for large \(x\)?

OpenStudy (anonymous):

Anyway well the answer is true because as we as x of this function f(x) tends to infinity the answer is +inf so its surely true

OpenStudy (anonymous):

@joyraheb that's actually very, very incorrect; sure, large is context-dependent, but once again the existence of the entire field of asymptotics disagrees with you

OpenStudy (anonymous):

i'm confused :/

OpenStudy (anonymous):

joyraheb's answer to the question and explanation are correct but his other comments are misleading

OpenStudy (anonymous):

You idiot we're talking about ordinary functions here who in the hell mention asymptotes?!

OpenStudy (usukidoll):

watch the language.

OpenStudy (anonymous):

Mentioned*

OpenStudy (anonymous):

Sorry 😂 i just get nervous quiqly

OpenStudy (anonymous):

note this is the same reason why the limit of rational functions works out as it does: $$\frac{ax^n+O(x^{n-1})}{bx^n+O(x^{n-1})}=\frac{a+O(1/x)}{b+O(1/x)}\sim\frac{a}b$$for large \(x\)

OpenStudy (alekos):

there are no asymptotes for this equation

OpenStudy (usukidoll):

Alright... but remember from OS Coc Be Nice - I will stay positive, be friendly, and not mean :)

OpenStudy (alekos):

its just a 3rd order polynomial

OpenStudy (anonymous):

@joyraheb @alekos I think you're both unfortunately confused

OpenStudy (anonymous):

okay... i'm really confused now?

OpenStudy (anonymous):

@alekos that's not what asymptotic analysis means

OpenStudy (usukidoll):

well... the only way to shed the light is to really find out what the question is really asking.

OpenStudy (anonymous):

https://en.wikipedia.org/wiki/Asymptotic_analysis

OpenStudy (anonymous):

I've already explained what the question is really asking -- it wants to know about end-behavior, which is essentially asking about asymptotic equivalence, i.e. knowing that \(x^3\) dominates as \(x\to\infty\), and since \(x^3\) 'rises' for large \(x\) then it follows that \(x^3-3x^2+2x\) will, too

OpenStudy (alekos):

there's no asymptotic line such as x=6 or y=-5. It just goes to plus infinity, plain and simple

OpenStudy (anonymous):

so it's true??

OpenStudy (anonymous):

@mishtea is it true for \(x^3\)? does \(x^3\) 'rise' for larger and larger \(x\)? is it increasing?

OpenStudy (anonymous):

@alekos you are still confused; please reread my previous comments and maybe check out the Wikipedia article on asymptotic analysis

OpenStudy (usukidoll):

maybe let x = any number and see if it increases ¯\_(ツ)_/¯

OpenStudy (anonymous):

yes i think because you said earlier that x^3 rises for the larger x and the equation that right

OpenStudy (usukidoll):

if the value is positive though... anyway I'm just going on the sidelines. This might turn ugly

OpenStudy (anonymous):

well, think of the graph? what happens as you move to the right? as you think of bigger and bigger \(x\)?

OpenStudy (anonymous):

it rises

OpenStudy (anonymous):

@mishtea have you studied limits, studying variations.... If not then you wouldn't understand because its a very simple question if you have taken these

OpenStudy (anonymous):

right, and since we've shown that for large \(x\), that since \(x^3\) grows so much faster than the \(-3x^2+2x\) part, we know that in the 'long run' the behavior of \(x^3-3x^2+2x\) will be like that of \(x^3\). so if you know that \(x^3\) will rise for large \(x\), then so will \(x^3-3x^2+2x\)

OpenStudy (anonymous):

... yes i have but i'm not a math person ...